Feldpausch Corporation has provided the following data from its activity-based costing system:

Activity Cost Pool Total Cost Total Activity
Assembly $ 1,150,100 62,000 machine-hours
Processing orders $ 54,554 1,860 orders
Inspection $ 194,310 2,550 inspection-hours

The company makes 990 units of product W26B a year, requiring a total of 1,610 machine-hours, 65 orders, and 30 inspection-hours per year. The product's direct materials cost is $52.35 per unit and its direct labor cost is $17.21 per unit. The product sells for $115.45 per unit.

According to the activity-based costing system, the product margin for product W26B is: (Round your intermediate calculations and final answers to 2 decimal places.)

(A) $11,373.15

(B) $45,431.10

(C) $13,659.15

(D) $13,279.60
Answer #1

Activity rate=Respective cost/Respective activity

Activity rate for:

Assembly=(1,150,100/62000)=$18.55 per machine hour

Processing orders=(54,554/1860)=$29.33 per order

Inspection=(194310/2550)=$76.2 per inspection hour

Hence total cost allocated=(18.55*1610)+(29.33*65)+(76.2*30)=$34057.95

Add:Direct materials=(990*52.35)=$51826.5

Add:Direct labor=(990*17.21)=$17037.9

Total cost=102922.35

Total revenue=(990*115.45)=$114295.5

Hence product margin=Total revenue-Total cost

=114295.5-102922.35

=$11373.15